Find an equation for the line parallel to y equals 4 x minus 6y=4x−6 with​ y-intercept left parenthesis 0 comma 3 right parenthesis(0,3). Write the answer in​ slope-intercept form.

Answer by Cromlix(4381)

Hi there,
Lines that are parallel to one another
have the same gradient.
m1 = m2
So m = 4
Using line equation:
y - b = m(x - a)
m = 4, (a, b) = (0, 3)
y - 3 = 4(x - 0)
y - 3 = 4x
y = 4x + 3
Hope this helps :-)